OpenStudy (anonymous):

Find the separable solutions to the differential equation (the heat equation) \[u_t=u_{xx}\] Given conditions: \[u_x(0,t)=0\]\[u_x(\pi,t)=0\] My professor didn't finish any example during class...

OpenStudy (anonymous):

how about initial condition?

OpenStudy (anonymous):

here is a complete solution of heat equation with L instead pi and with different boundary conditions i hope this may help u

OpenStudy (anonymous):

I'll read through this, thats all the info we're given

OpenStudy (anonymous):

Looks like parametric questions with t representing time. What do you mean by heat equation? I've quite familiar with heat ;-) Definition: Heat is the flow of thermal kinetic energy (temperature) from an region of high concentration to lower concentration as per natural entropy flow. The 3 modes of heat transfer are conduction, convection (which is really just advection plus conduction trough the no-slip zone), and radiation (transfer by electromagnetic particles) I think the document @mukushla gave you is probably your best bet.

OpenStudy (anonymous):

through* the no-slip zone

OpenStudy (anonymous):

Agent - It's a partial differential equation called the heat equation

OpenStudy (anonymous):

Context? FYI: heat \(\neq\) temperature In contrast to what most people say when they talk.

OpenStudy (anonymous):

I'm taking that this is from Differential Equation mathematics, a subject I've heard of but alas am not familiar with myself. My experience is from the experimental and application sides of heat. I guess what I was asking was if this "initial value problem" is in anyway related to a mode of heat transfer. I'll see if I can find some people who might be able to help with the mathematics explanation. :-)

OpenStudy (anonymous):

"The heat equation is an important partial differential equation which describes the distribution of heat (or variation in temperature) in a given region over time." - http://en.wikipedia.org/wiki/Heat_equation I can separate it to: \[\lambda = 0, X(x)=Ax+b\] \[\lambda <0, X(x)=Acosh(\sqrt{\mu}x)+Bsinh(\sqrt{\mu}x)\] \[\lambda >0,X(x)=Acos(\sqrt{\lambda}x)+Bsin(\sqrt{\lambda}x)\] Note sure exactly how to proceed, i'll read through the pdf.

OpenStudy (anonymous):

It's from Differential equations 2 class (partial differential equations, unlike diffeq 1 - ordinary differential equations.

OpenStudy (anonymous):

partial differential equation

OpenStudy (turingtest):

I'd have to review a lot of stuff to give you a full answer, so I'm just going to refer you to something similar (but perhaps a bit simpler) that what @mukushla gave you http://tutorial.math.lamar.edu/Classes/DE/SolvingHeatEquation.aspx

OpenStudy (turingtest):

again it has L instead of \(\pi\) but that shouldn't be a big problem to fix...

OpenStudy (anonymous):

Cool, i'll take a look through it

OpenStudy (anonymous):

Looks like i'm going to have to figure out eigenfunctions/values for systems of ode's, we skipped that chapter in ODE class...

OpenStudy (turingtest):

bummer... well all the info on how to do that is in those notes, but learning all that would probably take at least a day or two...

OpenStudy (anonymous):

In the case of \[\lambda =0\]\[X(x)=Ax+B\]\[X'(x)=A\]\[X'(0)=0=A \rightarrow A=0\] Thus,\[X(x)=B\]is a non-trivial solution In case of \[\lambda <0\]\[Let -\lambda = \mu\]\[X(x)=Acosh(\sqrt{\mu}x)+Bsinh(\sqrt{\mu}x)\]\[X'(x)=A \sqrt{\mu}\sinh(\sqrt{\mu}x)+B \sqrt{\mu}\cosh({\sqrt{\mu}x)}\]\[X'(0)=0=B \sqrt{\mu} \rightarrow B=0\]B must equal zero since lambda must be positive. \[X'(\pi)=0=A \sqrt{\mu}\sinh(\sqrt{\mu}\pi)\rightarrow A=0\]A must equal zero since mu <0 and pi is positive. No non-trivial solutions for lambda <0 In case of \[\lambda >0\]\[X(x)=Acos(\sqrt{\lambda}x)+Bsin(\sqrt{\lambda}x)\]\[X'(x)=-A \sqrt{\lambda}\sin(\sqrt{\lambda}x)+B \sqrt{\lambda}\cos(\sqrt{\lambda}x)\]\[X'(0)=0=B \sqrt{\lambda}\rightarrow B=0\]\[X'(\pi)=0=-A \sqrt{\lambda}\sin(\sqrt{\lambda}\pi)\]Here we have non trivial solutions when sqrt(lambda) is an integer, \[n=\sqrt{\lambda}\]Thus, \[X(x)=Acos(n \pi)\]Is a non-trivial solution when n=1,2,3,4... Continued on next post!

OpenStudy (anonymous):

Missed a posst, this one goes before the last one... From the last post we got the solutions for X(x) \[X_0(x)=B\]\[X_n(x)=Acos(nx)\]n = 1,2,3,4.... Now we need to find the corresponding solutions for T(t) \[T'(t)+\lambda T(t)=0\]\[T(t)=Ce^{-\lambda t}\] The corresponding solutions are: \[T_0(t)=C\]\[T_n(t)=Ce^{-n^2t}\]Note:\[\sqrt{\lambda}=n \rightarrow \lambda =n^2\] Since we assumed a solution of the form\[u(x,t)=X(x)T(t)\] \[u_0(x,t)=BC=E\]\[u_n(x,t)=De^{-n^2t}\cos(nx)\] These are the separable solutions, we must take into consideration the solutions that are a linear combination of these, next post!

OpenStudy (anonymous):

it deleted one of my posts... We were also given the initial condition:\[u(x,0)=f(x)\] We will put the solutions of the last post together in this way: \[u(x,t)=E+\sum_{n=1}^{\infty }Dcos(nx)\] I'll rewrite this in a convenient way that resembles a fourier series: \[u(x,t)=\frac{a_0}{2}+\sum_{n=1}^{\infty }a_n cos(nx)\] So we can now find a_n \[f(x)=u(x,0)=\frac{a_0}{2}+\sum_{n=1}^{\infty}a_ncos(nx)\] \[f(x)\cos(kx)=\frac{a_0}{2}\cos(kx)+\sum_{n=1}^{\infty}a_ncos(nx)\cos(kx)\] \[\int\limits_{0}^{\pi} f(x)\cos(kx)dx=\int\limits_{0}^{\pi} \frac{a_0}{2}\cos(kx)dx+\int\limits_{0}^{\pi} a_ncos(nx)\cos(kx)dx\] \[\int\limits_{0}^{\pi}f(x)\cos(nx)dx=a_n \int\limits_{0}^{\pi}\cos^2(nx)dx\] \[\int\limits_{0}^{\pi}f(x)\cos(nx)dx=\frac{a_n}{2}\int\limits_{0}^{\pi}[1+\cos(2nx)]dx\] \[\int\limits_{0}^{\pi}f(x)\cos(nx)dx=\frac{a_n \pi}{2}\] \[\frac{2}{\pi}\int\limits_{0}^{\pi}f(x)\cos(nx)dx=a_n\] Finally: \[u(x,t)=\frac{a_0}{2}+\sum_{n=1}^{\infty}\cos(nx)\int\limits_{0}^{\pi}\frac{2}{\pi}f(x)\cos(nx)dx\] Nasssty....

OpenStudy (anonymous):

thats the right order...